OpenStudy (anonymous):

f(x)= /x/ +10 ; f(-4) is it linear or not linear?

OpenStudy (shadowlegendx):

Hello Kirito-kun

OpenStudy (anonymous):

hello! do u think u can help meee? lol this is so hard for mee

OpenStudy (anonymous):

F(-4)= |-4| + 10 = 4+10=14 I think it's not linear but not so sure

OpenStudy (mayankdevnani):

@phuongduongm modulus can open with +ve value or with -ve value

OpenStudy (anonymous):

thanks guys for helping!! :D

OpenStudy (mayankdevnani):

@jellybean11122 here is your answer :- \[\large \bf f(x)=\left| x \right|+10\] Modulus can open with both +ve and -ve value. So,there are 2 cases. 1 Case :- when open with +ve value :- \[\large \bf f(-4)=+(-4)+10=-4+10=6......(1)Answer\] 2 Case :- when open with -ve value :- \[\large \bf f(-4)=-(-4)+10=4+10=14....(2)Answer\] So there are 2 possible answers. And both answers have not any variable they all are constant. So,they are NOT LINEARS.

OpenStudy (mayankdevnani):

understood? @jellybean11122

OpenStudy (anonymous):

thanks so much i understand now! thx thx thx :)) @mayankdevnani

OpenStudy (mayankdevnani):

welcome :)